Show that whether x5 + 10x3 + x + 1 is O(x4) or not?


1
Expert's answer
2021-10-21T10:37:34-0400

"\dfrac{x^5+10x^3+x+1}{x^4}= \dfrac{x^5}{x^4}+\dfrac{10x^3}{x^4}+\dfrac{x}{x^4}+\dfrac{1}{x^4}=x+\dfrac{10}{x}+\dfrac{1}{x^3}+\dfrac{1}{x^4}"


"\implies x+10x^{-1}+x^{-3}+x^{-4}"


Degree of polynomial = 1

So, it is not "O(x^4)"